OpenStudy (anonymous):

Simplify. −1w + 4w4 + y4 – 3y4 – 2w – 3w4 A. −2w4 – 2y4 B. −w4y4 – 3w C. w4 – 3w – 2y4 D. 12w4 + 2w – 3y4

OpenStudy (anonymous):

lets put our like terms together.. -1w - 2w = -3w 4w^4 - 3w^4 = w^4 y^4 - 3y^4 = - 2y^4 now put it all together.. w^4 - 3w - 2y^4 any questions ?

OpenStudy (anonymous):

like terms will have the same variable and the same exponent. If they are not like terms, you cannot add/subtract.

OpenStudy (anonymous):

thanks kellie